# The underlying structures of self-attention: symmetry, directionality, and emergent dynamics in Transformer training

### v1

@@ Acknowledgments
To demonstrate the column dominance, we first show that if one probability distribution has a higher variance than another, it is more likely to produce samples with higher values. We formalize this in the following Lemma.
@@
Let two probability distributions $a,b$ have the same mean $\mu$ and different variance $\sigma_{a}^{2}>\sigma_{b}^{2}$. The probability of a random sample of each distribution being larger than a value $z$ is higher for samples from $a$ than from $b$ for $z>\sqrt{\sigma_{a}\sigma_{b}}-\mu$,
$\text{Pr}\left[X_{a}>z\right]\geq\text{Pr}\left[X_{b}>z\right]\quad\forall z>\sqrt{\sigma_{a}\sigma_{b}}-\mu$ (S88)
@@
We now formalize the connection between this result and the asymmetric growth of columns in the following Theorem.
Theorem S1.12.
(Autoregressive training induces directionality) Let $\mathcal{V}=[t_{0},\dots,t_{V}]$ be a set of tokens. Let $U$ be an ordered sequence of $N$ tokens $U=[t_{1},\dots,t_{N}]$ where $\Pr[t_{j}=t^{*}]$ is the probability that the token at index $j$ in $U$ is given by $t^{*}\in\mathcal{V}$. Let $\{\mathbf{x}_{1}\dots,\mathbf{x}_{n}\}$ be the embedding associated with $\mathcal{V}$ such that each $\mathbf{x}_{i}\sim\mathcal{D}$ is drawn i.i.d. from a probability distribution $\mathcal{D}$ with zero mean and semipositive covariance $\text{Cov}(\mathbf{x}_{i})\neq\sigma_{x}^{2}\mathbb{I}$. Let $\Delta\mathbf{W}_{qk}$ be the weight update of $\mathbf{W}_{qk}$ from Proposition 2.2. Let $\Delta\mathbf{W}_{qk}$ be derived from an autoregressive objective function as in Definition S1.3. Then, there exists a value $\gamma\in\mathbb{R}$ such that,
$\text{Pr}[||\mathbf{w}_{\cdot,k}||>w]>\text{Pr}[||\mathbf{w}_{m,\cdot}||>w]\,\,\ \forall\,w>\gamma$ (S94)
@@
Let $\mathbb{E}[\mu_{c}(t^{*})]$ be the expected number of tokens that are predicted by a given token $t^{*}$, and $\mathbb{E}[\mu_{p}(t^{*})]$ the expected number of tokens that predict a given token $t^{*}$. When assuming autoregressive training, it follows from Proposition S1.10 that
$\frac{\mathbb{E}[\mu_{c}(t^{*})]}{\mathbb{E}[\mu_{p}(t^{*})]}=\frac{\sum_{j=1}^{N}(N-j)\Pr[t_{j}=t^{*}]}{\sum_{j=1}^{N}(j-1)\Pr[t_{j}=t^{*}]}\,.$ (S95)
Let the statistical correlation between tokens in $U$,
$\frac{\mathbb{E}[\mu_{c}(t^{*})]}{\mathbb{E}[\mu_{p}(t^{*})]}>1\,,$ (S96)
for most $t^{*}\in\mathcal{V}$. It follows that the average number of tokens that are predicted by $t^{*}$ is smaller than the average number of tokens that predict it. Therefore, the number of times $t^{*}$ contributes to the weight update of $\mathbf{W}_{qk}$ as predicted token is on average smaller than the number of times it contributes as context token. It follows from Proposition S1.8 that the number of times the following is true,
@@
is on average higher than the number of times the following is true,
$\frac{\mathbb{E}\left[||\Delta\mathbf{w}_{\cdot,k}||^{2}\right]}{\mathbb{E}\left[||\Delta\mathbf{w}_{m,\cdot}||^{2}\right]}<1\,.$ (S98)
Therefore, the net increase of the norm of the columns is bigger than the net increase of the norm of the rows. We now prove that the number of columns with a high norm exceeds the number of rows with a high norm. At initialization, let each $(k,m)$ entry of $\mathbf{W}_{qk}$ be an i.i.d. random variable drawn from a probability distribution with mean $\mu$ and variance $\sigma^{2}$. It follows that the mean norm of rows and columns are equal at initialization, as follows,
$\mathbb{E}\left[|\mathbf{w}_{m,\cdot}|\right]\bigg{|}_{\text{init}}=\mathbb{E}\left[|\mathbf{w}_{\cdot,k}|\right]\bigg{|}_{\text{init}}=\mu\,\,\forall k,m\,.$ (S99)
Furthermore, the variance of the norms of the rows $\sigma_{\text{rows}}^{2}$ and of the columns $\sigma_{\text{cols}}^{2}$ are equal at initialization,
$\text{Var}\left(|\mathbf{w}_{m,\cdot}|\right)\bigg{|}_{\text{init}}=\text{Var}\left(|\mathbf{w}_{\cdot,k,\cdot}|\right)\bigg{|}_{\text{init}}=\sigma^{2}\,\,\forall k,m\,.$ (S100)
Therefore, the following inequality holds at the end of training,
$\text{Var}\left(|\mathbf{w}_{m,\cdot}|\right)>\text{Var}\left(|\mathbf{w}_{\cdot,k}|\right)\,\,\forall k,m\,.$ (S101)
Finally, we obtain from Lemma S1.11 that the following inequality holds,
$\text{Pr}[||\mathbf{w}_{\cdot,k}||>w]>\text{Pr}[||\mathbf{w}_{m,\cdot}||>w]\,\,\ \forall\,w>\gamma$ (S102)
where $\gamma=\sqrt{\text{Var}\left(|\mathbf{w}_{m,\cdot}|\right)\text{Var}\left(|\mathbf{w}_{\cdot,k}|\right)}-\mu$, thus concluding the proof. ∎
S1.4.5 Theorem on the emergence of symmetry
@@
(Bidirectional training induces rows and columns with equal norm) Let $\mathcal{V}=[t_{0},\dots,t_{V}]$ be a set of tokens. Let $U$ be an ordered sequence of $N$ tokens $U=[t_{1},\dots,t_{N}]$ where $\Pr[t_{j}=t^{*}]$ is the probability that the token at index $j$ in $U$ is given by $t^{*}\in\mathcal{V}$. Let $\{\mathbf{x}_{1}\dots,\mathbf{x}_{n}\}$ be the random embedding of the tokens as in Theorem S1.12. Let $\Delta\mathbf{W}_{qk}$ be the weight update of $\mathbf{W}_{qk}$ from Proposition 2.2. Let $\Delta\mathbf{W}_{qk}$ be derived from a bidirectional objective function as in Definition S1.3. Then, the variance of the norm of the rows and the columns at the end of training is equal,
$\text{Var}\left(|\mathbf{w}_{m,\cdot}|\right)=\text{Var}\left(|\mathbf{w}_{\cdot,k}|\right)\,\,\forall k,m\,.$ (S103)
@@
It follows from Proposition S1.10 that the number of times the following is true,
is on average equal to the number of times the following is true,
$\frac{\mathbb{E}\left[||\Delta\mathbf{w}_{\cdot,k}||^{2}\right]}{\mathbb{E}\left[||\Delta\mathbf{w}_{m,\cdot}||^{2}\right]}<1\,.$ (S105)
Therefore, the net increase of the norm of the columns is equal to the net increase of the norm of the rows. Following the same initialization assumptions as in Theorem S1.12, it follows that at the end of training,
$\text{Var}\left(|\mathbf{w}_{m,\cdot}|\right)=\text{Var}\left(|\mathbf{w}_{\cdot,k}|\right)\,\,\forall k,m\,.$ (S106)
thus concluding the proof. ∎
Now, we show that the bidirectional nature of the training objective is such that every pair of tokens $(i,j)$ in a given sequence contributes to a term that is approximately symmetric. Here, we assume that predicting the $i$-th token from the $j$-th token is correlated to predicting the $j$-th token from the $i$-th token. In other words, predicting one token from the other gives similar predictions, that is, the term $\beta_{ij}$ and $\beta_{ji}$ from Proposition 2.2 are correlated. We formalize this in the following Theorem.
Theorem S1.14.
(Bidirectional training induces symmetry) Let $U$ be an ordered sequence of $N$ tokens $U=[t_{1},\dots,t_{N}]$. Let $\Delta\mathbf{W}_{qk}$ be derived from a bidirectional objective function as in Definition S1.3. It follows that the weight update of $\mathbf{W}_{qk}^{l}$ is given by
$\Delta\mathbf{W}_{qk}^{l}=\sum_{i=1}^{N}\sum_{j=1}^{N}\beta^{l}_{ij}\mathbf{K}^{l-1}_{ij}\,,$ (S107)
$\Delta\mathbf{W}_{qk}^{l}\big{|}_{\mathbf{t}_{i}\leftrightarrow\mathbf{t}_{j}}=\beta_{ij}^{l}\mathbf{K}^{l-1}_{ij}+\beta_{ji}^{l}{\mathbf{K}^{l-1}_{ij}}^{T}\,,$ (S108)
that is approximately symmetric,
$\Delta\mathbf{W}_{qk}^{l}\big{|}_{\mathbf{t}_{i}\leftrightarrow\mathbf{t}_{j}}\approx\Delta\mathbf{W}_{qk}^{l}\big{|}^{\top}_{\mathbf{t}_{i}\leftrightarrow\mathbf{t}_{j}}\,.$ (S109)
@@
Let $U$ be an ordered sequence of $N$ tokens $U=[t_{1},\dots,t_{N}]$. It follows from Proposition 2.2 that the weight update for $\mathbf{W}_{qk}^{l}$ following the gradient of $\mathcal{L}(\mathbf{U})$ w.r.t. $\mathbf{W}_{qk}^{l}$ is given by
$\Delta\mathbf{W}_{qk}^{l}=\sum_{i=1}^{N}\sum_{j=1}^{N}\beta^{l}_{ij}\mathbf{K}^{l-1}_{ij}\,,$ (S110)
where we neglect any constant of proportionality - such as a learning rate - for simplicity. The double summation in Equation (S110) contains $N^{2}$ elements. We can rewrite the double summation as follows,
$\sum_{i=1}^{N}\sum_{j=1}^{N}\beta^{l}_{ij}\mathbf{K}^{l-1}_{ij}=\sum_{i=1}^{N}\beta^{l}_{ii}\mathbf{K}^{l-1}_{ii}+\sum_{\begin{subarray}{c}i,j=1\\ i<j\end{subarray}}^{N}(\beta^{l}_{ij}\mathbf{K}^{l-1}_{ij}+\beta^{l}_{ji}\mathbf{K}^{l-1}_{ji})$ (S111)
@@
$\sum_{\begin{subarray}{c}i,j=1\\ i<j\end{subarray}}^{N}(\beta^{l}_{ij}\mathbf{K}^{l-1}_{ij}+\beta^{l}_{ji}\mathbf{K}^{l-1}_{ji})=\sum_{\begin{subarray}{c}i,j=1\\ i<j\end{subarray}}^{N}(\beta^{l}_{ij}\mathbf{K}^{l-1}_{ij}+\beta^{l}_{ji}{\mathbf{K}^{l-1}_{ij}}^{\top})\,,$ (S112)
$\begin{split}\sum_{\begin{subarray}{c}i,j=1\\ i<j\end{subarray}}^{N}(\beta^{l}_{ij}\mathbf{K}^{l-1}_{ij}+\beta^{l}_{ji}{\mathbf{K}^{l-1}_{ij}}^{\top})&=\sum_{\begin{subarray}{c}i,j=1\\ i<j\end{subarray}}^{N}\big{[}\beta_{ij}^{l}(\mathbf{S}^{l-1}_{ij}+\mathbf{N}^{l-1}_{ij})+\beta_{ji}^{l}(\mathbf{S}^{l-1}_{ij}+\mathbf{N}^{l-1}_{ij})^{T}\big{]}\\ &=\sum_{\begin{subarray}{c}i,j=1\\ i<j\end{subarray}}^{N}\big{[}(\beta^{l}_{ij}+\beta^{l}_{ji})\mathbf{S}^{l-1}_{ij}+(\beta^{l}_{ij}-\beta^{l}_{ji})\mathbf{N}^{l-1}_{ij}\big{]}\,.\end{split}$ (S113)
Let $\Delta\mathbf{W}_{qk}^{l}\big{|}_{\mathbf{t}_{i}\leftrightarrow\mathbf{t}_{j}}=\beta_{ij}^{l}\mathbf{K}^{l-1}_{ij}+\beta_{ji}^{l}{\mathbf{K}^{l-1}_{ij}}^{T}$, and let $\beta^{l}_{ij}$ and $\beta^{l}_{ji}$ be such that $\text{sign}(\beta^{l}_{ij})=\text{sign}(\beta^{l}_{ji})$ and $|\beta^{l}_{ij}|\approx$$|\beta^{l}_{ji}|$. It follows that
$\Delta\mathbf{W}_{qk}^{l}\big{|}_{\mathbf{t}_{i}\leftrightarrow\mathbf{t}_{j}}\approx\sum_{\begin{subarray}{c}i,j=1\\ i<j\end{subarray}}^{N}\beta^{l}_{ij}\mathbf{S}^{l-1}_{ij}=\sum_{\begin{subarray}{c}i,j=1\\ i<j\end{subarray}}^{N}\beta^{l}_{ij}{\mathbf{S}^{l-1}}^{\top}_{ij}=\Delta{\mathbf{W}_{qk}^{l}}^{\top}\big{|}_{\mathbf{t}_{i}\leftrightarrow\mathbf{t}_{j}}\,$ (S114)
@@
S1.5. Properties of the symmetry score in Definition 3.1 and related proofs
@@
S1.6. Properties of the directionality score in Definition 3.2 and related proofs
@@
We trained two BERT models [Devlin et al., 2019] to examine the evolution of the symmetry and directionality scores throughout the training process. Detailed information regarding the training procedure is provided below.
@@
Configuration BERT BERT-Mini
Hidden Size 768 256
Intermediate Size 3072 1024
Number of Attention Heads 12 4
Number of Hidden Layers 12 4
Attention Dropout Probability 0.1 0.1
Hidden Activation Function gelu gelu
Hidden Dropout Probability 0.1 0.1
Layer Normalization Epsilon 1e-12 1e-12
Max Position Embeddings 512 512
Position Embedding Type absolute absolute
Vocabulary Size 30522 30522
Table S1: Configurations for BERT and BERT-Mini models.
We train the standard BERT model (referred to as BERT) and a smaller version (referred to as BERT-Mini), following the implementation by [Devlin et al., 2019]. Table S1 provides an overview of the model parameters. The standard BERT model has $12$ layers, $12$ attention heads, and embedding dimensions of $768$ for the hidden layers and $3072$ for the intermediate layers. In contrast, the smaller BERT-Mini model uses $4$ layers, $4$ attention heads, and embedding dimensions of $256$ and $1024$, respectively.
@@
We optionally initialize the models with either symmetric or skew-symmetric attention weights by initializing the key and query weight matrices of all self-attention blocks following a specific procedure. For symmetric initialization, the query weight matrix $\mathbf{W_{q}}$ is initialized randomly, and then the key weight matrix is set to $\mathbf{W_{k}}=\mathbf{W_{q}}$. For skew-symmetric initialization, $\mathbf{W_{q}}$ is initialized randomly, and a skew-symmetric matrix $\mathbf{S}$ is generated by first initializing a random matrix, $\mathbf{A}$ and then calculating $\mathbf{S}=\mathbf{A}-\mathbf{A}^{\top}$. The key matrix is then defined as $\mathbf{W_{k}}=\mathbf{W_{q}}\mathbf{S}^{\top}$, ensuring that the matrix multiplication $\mathbf{W_{q}}\mathbf{W_{k}}^{\top}$ is skew-symmetric.
Loss Constraints
The loss function $\mathcal{L}(\mathbf{U})$ described in Equation (S22) is optionally extended with a constraint to either encourage symmetric or skew-symmetric attention weights.
$\mathcal{L}(\mathbf{U}):=\mathcal{L}(\mathbf{U})+\gamma\mathcal{L}_{c}(\mathcal{W}).$ (S146)
Here, $\gamma$ is a coefficient defining the importance of the $\mathcal{L}_{c}$ constraint. To encourage symmetry, we minimize the condition
$\mathcal{L}_{c}(\mathcal{W})=\frac{\mathbf{M}^{2}}{\mathbf{M}_{s}^{2}}.$ (S147)
This constraint reaches its minimum if all attention weights have a symmetry score $s=1$. To encourage skew-symmetry, we minimize the condition
$\mathcal{L}_{c}(\mathcal{W})=\frac{\mathbf{M}^{2}}{\mathbf{M}_{n}^{2}},$ (S148)
leading to a symmetry score of $s=0$.

### v2

@@ Acknowledgments
Before proving column dominance in the next Theorem, we first establish that a probability distribution with higher variance is more likely to produce large values. Specifically, when two distributions share the same mean but differ in variance, the one with higher variance has a greater probability of generating samples that exceed a given threshold. We formalize this in the following lemma.
@@
Let $a$ and $b$ be two probability distributions with the same mean $\mu$, but different variances such that $\sigma_{a}^{2}>\sigma_{b}^{2}$. Then, for all values $z>\sqrt{\sigma_{a}\sigma_{b}}-\mu$, the probability that a random sample from distribution $a$ exceeds $z$ is greater than or equal to the probability that a sample from distribution $b$ exceeds $z$, that is,
$\Pr[X_{a}>z]\geq\Pr[X_{b}>z]\quad\text{for all }z>\sqrt{\sigma_{a}\sigma_{b}}-\mu\,.$ (S83)
@@
S1.5.5 Proof of Theorem 2.3
Here, we show that the weight updates of $\mathbf{W}_{qk}$ induce column dominance during autoregressive training, leading to the emergence of directionality. To do so, we make the following assumptions:
  • We adopt the same assumptions stated in Propositions S1.7 and S1.9.
  • At initialization, the entries of $\mathbf{W}_{qk}$ are drawn from a probability distribution $\mathcal{P}$ with finite mean $\mu$ and variance $\sigma^{2}$. This is satisfied by any standard machine learning initialization scheme. This second assumption implies that each $k$-th column $\mathbf{w}_{\cdot,k}$ and $m$-th row $\mathbf{w}_{m,\cdot}$ have the same mean $\mu$ and variance $\sigma^{2}/\sqrt{n}$. Furthermore, the squared norm of both rows and columns has equal mean $n(\sigma^{2}+\mu^{2})$ and variance $n(2\sigma^{4}+4\mu^{2})$.
@@
Let $\{U_{1},\dots,U_{D}\}$ be a dataset of sequences of length $N$ drawn i.i.d. from a distribution $P$. Let $k_{u}$ denote the position of token $t^{*}$ in sequence $U_{u}$, and let superscript $u$ refer to tokens in the $u$-th sequence. Following Proposition 2.2 and S1.7, we obtain
$\Delta\mathbf{w}_{\cdot,k}=\sum_{u=1}^{D}\left[[\mathbf{y}]_{k}\sum_{i=k_{u}+1}^{N}\beta^{u}_{i*}\mathbf{x}^{u}_{i}+\sum_{j=1}^{k_{u}-1}\beta^{u}_{*j}[\mathbf{x}^{u}_{j}]_{k}\mathbf{y}\right]\,,$ (S89)
while the squared norm of the weight update of the $m$-th row by
$\Delta\mathbf{w}_{m,\cdot}=\sum_{u=1}^{D}\left[\sum_{i=k_{u}+1}^{N}\beta^{u}_{i*}[\mathbf{x}^{u}_{i}]_{m}\mathbf{y}+[\mathbf{y}]_{m}\sum_{j=1}^{k_{u}-1}\beta^{u}_{*j}\mathbf{x}^{u}_{j}\right]\,,$ (S90)
where in both equations the superscript $u$ indicates tokens belonging to the $u$-th sequence in the dataset, and $k_{u}$ indicates the position index of $t^{*}$. It follows again from Proposition S1.7 that the expected value of the squared norm of these updates is
$\frac{\mathbb{E}_{\mathcal{D}}\left[||\Delta\mathbf{w}_{\cdot,k}||^{2}\right]}{\mathbb{E}_{\mathcal{D}}\left[||\Delta\mathbf{w}_{m,\cdot}||^{2}\right]}\approx\frac{\Gamma_{kk}\text{Tr}(\Sigma)\big{(}\sum_{u=1}^{D}\sum_{i=k_{u}+1}^{N}\beta^{u}_{i*}\big{)}+\Sigma_{kk}\text{Tr}(\Gamma)\big{(}\sum_{u=1}^{D}\sum_{j=1}^{k_{u}-1}\beta^{u}_{*j}\big{)}}{\Sigma_{mm}\text{Tr}(\Gamma)\big{(}\sum_{u=1}^{D}\sum_{i=k_{u}+1}^{N}\beta^{u}_{i*}\big{)}+\Gamma_{mm}\text{Tr}(\Sigma)\big{(}\sum_{u=1}^{D}\sum_{j=1}^{k_{u}-1}\beta^{u}_{*j}\big{)}}\,.$ (S91)
Following Proposition S1.9, let $\mu_{c}(t^{*},U)$ be the random variable quantifying the number of tokens predicted by $t^{*}$, while $\mu_{p}(t^{*},U)$ is the random variable quantifying the number of tokens predicted by $t^{*}$. When assuming autoregressive training, it follows that the empirical average of $\mu_{c}(t^{*},U)$ and $\mu_{p}(t^{*},U)$ over the dataset $\{U_{1},\dots,U_{D}\}$ approaches,
$\frac{\mathbb{E}_{P}[\mu_{c}(t^{*},U)]}{\mathbb{E}_{P}[\mu_{p}(t^{*},U)]}=\frac{\sum_{j=1}^{N}(N-j)\Pr[t_{j}=t^{*}]}{\sum_{j=1}^{N}(j-1)\Pr[t_{j}=t^{*}]}\,.$ (S92)
Following Remark S1.10, assume that the expected position $\mu(t^{*})$ of $t^{*}$ satisfies $\mu(t^{*})<(N+1)/2-\delta$ for some $\delta>0$, then
$\frac{\mathbb{E}_{P}[\mu_{c}(t^{*},U)]}{\mathbb{E}_{P}[\mu_{p}(t^{*},U)]}>1+\epsilon(t^{*})\,,$ (S93)
for some $\epsilon(t^{*})>0$. Hence, across the dataset, $t^{*}$ appears more often as a context token than as a predicted token. This implies $\mathbb{E}_{P}[k_{u}]<(N+1)/2$, so that the first terms in the numerator and denominator of Equation (S91) dominate, leading to,
@@
that is, the net increase in column norms exceeds that of row norms. Assume $\mathbf{W}_{qk}$ is initialized with i.i.d. entries from a distribution with mean $\mu$ and variance $\sigma^{2}$, then at the beginning of training $\mathrm{Var}(\|\mathbf{w}_{\cdot,k}\|)=\mathrm{Var}(\|\mathbf{w}_{m,\cdot}\|)$, while after training we have $\mathrm{Var}(\|\mathbf{w}_{\cdot,k}\|)>\mathrm{Var}(\|\mathbf{w}_{m,\cdot}\|)\,\,\forall k,m.$ Applying Lemma S1.11, it follows that for all $w>\gamma$, where $\gamma=\sqrt{\mathrm{Var}(\|\mathbf{w}_{\cdot,k}\|)\cdot\mathrm{Var}(\|\mathbf{w}_{m,\cdot}\|)}-\mu$,
$\text{Pr}[||\mathbf{w}_{\cdot,k}||>w]>\text{Pr}[||\mathbf{w}_{m,\cdot}||>w]\,\,\ \forall\,w>\gamma\,,$ (S95)
thus concluding the proof. ∎
S1.5.6 Proof of Theorem 2.4 and related remarks
@@
(Bidirectional training does not induce directionality) Let $\mathcal{V}=\{t_{0},\dots,t_{V}\}$ be a vocabulary of tokens, and let $\mathcal{U}\subset\mathcal{V}^{N}$ denote the sample space of all sequences of length $N$. Let $U=\{t_{1},\dots,t_{N}\}\in\mathcal{U}$ be a random variable with $U\sim P(U)$. Define $\Pr[t_{j}=t^{*}]$ as the marginal probability that the token at position $j$ equals $t^{*}\in\mathcal{V}$. Let $\{\mathbf{x}_{1},\dots,\mathbf{x}_{N}\}$ be the token embeddings corresponding to the elements of $U$, where each embedding $\mathbf{x}_{i}\sim\mathcal{D}$ is drawn i.i.d. from a distribution $\mathcal{D}$ with zero mean and covariance matrix $\text{Cov}(\mathbf{x}_{i})=\Sigma$.. Let $\mathbf{W}_{qk}$ be query-key matrix of a self-attention mechanism, and let $\Delta\mathbf{W}_{qk}$ denote its gradient update as defined in Proposition 2.1, computed under an autoregressive objective as in Definition S1.3. Then, the variance of the norm of the rows and the columns at the end of training is equal,
$\text{Var}\left(||\mathbf{w}_{m,\cdot}||\right)=\text{Var}\left(||\mathbf{w}_{\cdot,k}||\right)\,\,\forall k,m\,.$ (S96)
@@
Let $\{U_{1},\dots,U_{D}\}$ be a dataset of sequences of length $N$ drawn i.i.d. from a distribution $P$. Let $k_{u}$ denote the position of token $t^{*}$ in sequence $U_{u}$, and let superscript $u$ refer to tokens in the $u$-th sequence. Following Proposition 2.2 and S1.7, we obtain
$\Delta\mathbf{w}_{\cdot,k}=\sum_{u=1}^{D}\left[[\mathbf{y}]_{k}\sum_{i=k_{u}+1}^{N}\beta^{u}_{i*}\mathbf{x}^{u}_{i}+\sum_{j=1}^{k_{u}-1}\beta^{u}_{*j}[\mathbf{x}^{u}_{j}]_{k}\mathbf{y}\right]\,,$ (S97)
while the squared norm of the weight update of the $m$-th row by
$\Delta\mathbf{w}_{m,\cdot}=\sum_{u=1}^{D}\left[\sum_{i=k_{u}+1}^{N}\beta^{u}_{i*}[\mathbf{x}^{u}_{i}]_{m}\mathbf{y}+[\mathbf{y}]_{m}\sum_{j=1}^{k_{u}-1}\beta^{u}_{*j}\mathbf{x}^{u}_{j}\right]\,,$ (S98)
where in both equations the superscript $u$ indicates tokens belonging to the $u$-th sequence in the dataset, and $k_{u}$ indicates the position index of $t^{*}$. It follows again from Proposition S1.7 that the expected value of the squared norm of these updates is
$\frac{\mathbb{E}_{\mathcal{D}}\left[||\Delta\mathbf{w}_{\cdot,k}||^{2}\right]}{\mathbb{E}_{\mathcal{D}}\left[||\Delta\mathbf{w}_{m,\cdot}||^{2}\right]}\approx\frac{\Gamma_{kk}\text{Tr}(\Sigma)\big{(}\sum_{u=1}^{D}\sum_{i=k_{u}+1}^{N}\beta^{u}_{i*}\big{)}+\Sigma_{kk}\text{Tr}(\Gamma)\big{(}\sum_{u=1}^{D}\sum_{j=1}^{k_{u}-1}\beta^{u}_{*j}\big{)}}{\Sigma_{mm}\text{Tr}(\Gamma)\big{(}\sum_{u=1}^{D}\sum_{i=k_{u}+1}^{N}\beta^{u}_{i*}\big{)}+\Gamma_{mm}\text{Tr}(\Sigma)\big{(}\sum_{u=1}^{D}\sum_{j=1}^{k_{u}-1}\beta^{u}_{*j}\big{)}}\,.$ (S99)
Following Proposition S1.9, let $\mu_{c}(t^{*},U)$ be the random variable quantifying the number of tokens predicted by $t^{*}$, while $\mu_{p}(t^{*},U)$ is the random variable quantifying the number of tokens predicted by $t^{*}$. When assuming bidirectional training training, it follows that the empirical average of $\mu_{c}(t^{*},U)$ and $\mu_{p}(t^{*},U)$ over the dataset $\{U_{1},\dots,U_{D}\}$ approaches,
$\frac{\mathbb{E}_{P}[\mu_{c}(t^{*},U)]}{\mathbb{E}_{P}[\mu_{p}(t^{*},U)]}=1\,.$ (S100)
From Proposition S1.7, it follows that the net increase in the norms of the columns and rows of $\mathbf{W}_{qk}$ is the same. Assuming the same initialization conditions as in Theorem 2.3, it follows that by the end of training,
$\mathrm{Var}(\|\mathbf{w}_{m,\cdot}\|)=\mathrm{Var}(\|\mathbf{w}_{\cdot,k}\|)\quad\forall\,k,m\,,$ (S101)
which concludes the proof. ∎
We now prove Theorem 2.4, showing that the bidirectional training objective induces approximate symmetry in the gradient contributions of each token pair $(i,j)$. Specifically, we assume that predicting token $t_{i}$ from $t_{j}$ is correlated with predicting $t_{j}$ from $t_{i}$, that is, the terms $\beta_{ij}$ and $\beta_{ji}$ from Proposition 2.1 are correlated, leading to similar contributions in both directions.
@@
Let $U=\{t_{1},\dots,t_{N}\}$ be a sequence of tokens. It follows from Proposition 2.1 that the implciit weight update for $\mathbf{W}_{qk}$ following the gradient of $\mathcal{L}(\mathbf{U})$ w.r.t. $\mathbf{W}_{qk}$ is given by
$\Delta\mathbf{W}_{qk}=\sum_{i=1}^{N}\sum_{j=1}^{N}\beta_{ij}\mathbf{K}_{ij}\,,$ (S102)
where we neglect any constant of proportionality (e.g. learning rate) for simplicity. The double summation in Equation (S102) contains $N^{2}$ elements. Note that $\mathbf{K}_{ji}=\mathbf{K}_{ij}^{\top}$, so we can rewrite the double summation as follows,
$\sum_{i=1}^{N}\sum_{j=1}^{N}\beta_{ij}\mathbf{K}_{ij}=\sum_{i=1}^{N}\beta_{ii}\mathbf{K}_{ii}+\sum_{\begin{subarray}{c}i,j=1\\ i<j\end{subarray}}^{N}(\beta_{ij}\mathbf{K}_{ij}+\beta_{ji}\mathbf{K}_{ji})$ (S103)
@@
$\sum_{\begin{subarray}{c}i,j=1\\ i<j\end{subarray}}^{N}(\beta_{ij}\mathbf{K}_{ij}+\beta_{ji}\mathbf{K}_{ji})=\sum_{\begin{subarray}{c}i,j=1\\ i<j\end{subarray}}^{N}(\beta_{ij}\mathbf{K}_{ij}+\beta_{ji}{\mathbf{K}_{ij}}^{\top})\,,$ (S104)
and by decomposing $\mathbf{K}_{ij}$ in its symmetric and skew-symmetric parts, such that $\mathbf{K}_{ij}=\mathbf{S}_{ij}+\mathbf{N}_{ij}$, we obtain,
$\begin{split}\sum_{\begin{subarray}{c}i,j=1\\ i<j\end{subarray}}^{N}(\beta_{ij}\mathbf{K}_{ij}+\beta_{ji}{\mathbf{K}_{ij}}^{\top})&=\sum_{\begin{subarray}{c}i,j=1\\ i<j\end{subarray}}^{N}\big{[}\beta_{ij}(\mathbf{S}_{ij}+\mathbf{N}_{ij})+\beta_{ji}(\mathbf{S}_{ij}+\mathbf{N}_{ij})^{T}\big{]}\\ &=\sum_{\begin{subarray}{c}i,j=1\\ i<j\end{subarray}}^{N}\big{[}(\beta_{ij}+\beta_{ji})\mathbf{S}_{ij}+(\beta_{ij}-\beta_{ji})\mathbf{N}_{ij}\big{]}\,.\end{split}$ (S105)
Let $\beta_{ij}$ and $\beta_{ji}$ be such that $\text{sign}(\beta_{ij})=\text{sign}(\beta_{ji})$ and $|\beta_{ij}|\approx$$|\beta_{ji}|$. By defining $\Delta\mathbf{W}_{qk}\big{|}_{\mathbf{t}_{i}\leftrightarrow\mathbf{t}_{j}}=\beta_{ij}\mathbf{K}_{ij}+\beta_{ji}{\mathbf{K}_{ij}}^{T}$, it follows that
$\Delta\mathbf{W}_{qk}\big{|}_{\mathbf{t}_{i}\leftrightarrow\mathbf{t}_{j}}\approx\sum_{\begin{subarray}{c}i,j=1\\ i<j\end{subarray}}^{N}\beta_{ij}\mathbf{S}_{ij}=\sum_{\begin{subarray}{c}i,j=1\\ i<j\end{subarray}}^{N}\beta_{ij}{\mathbf{S}}^{\top}_{ij}=\Delta{\mathbf{W}_{qk}}^{\top}\big{|}_{\mathbf{t}_{i}\leftrightarrow\mathbf{t}_{j}}\,$ (S106)
@@
S1.6. Properties of the symmetry score in Definition 3.1 and related proofs
@@
S1.7. Properties of the directionality score in Definition 3.2 and related proofs
@@
We trained three BERT models [Devlin et al., 2019] to examine the evolution of the symmetry and directionality scores throughout the training process. Detailed information regarding the training procedure is provided below.
@@
Configuration BERT BERT-Mini BERT-Large
Hidden Size 768 256 1024
Intermediate Size 3072 1024 4096
Number of Attention Heads 12 4 16
Number of Hidden Layers 12 4 24
Attention Dropout Probability 0.1 0.1 0.1
Hidden Activation Function gelu gelu gelu
Hidden Dropout Probability 0.1 0.1 0.1
Layer Normalization Epsilon 1e-12 1e-12 1e-12
Max Position Embeddings 512 512 512
Position Embedding Type absolute absolute absolute
Vocabulary Size 30522 30522 30522
Table S1: Configurations for BERT, BERT-Mini, and BERT-Large models.
We train the standard BERT model (referred to as BERT), a smaller version (referred to as BERT-Mini), and a larger version (referred to as BERT-Large), following the implementation by [Devlin et al., 2019]. Table S1 provides an overview of the model parameters. The standard BERT model has $12$ layers, $12$ attention heads, and embedding dimensions of $768$ for the hidden layers and $3072$ for the intermediate layers. In contrast, the smaller BERT-Mini model uses $4$ layers, $4$ attention heads, and embedding dimensions of $256$ and $1024$, respectively. The larger BERT-Large model features $24$ layers, $16$ attention heads, and embedding dimensions of $1024$ and $4096$ for the hidden and intermediate layers, respectively.
@@
We optionally initialize the models using symmetric attention weights. Specifically, in each self-attention block, the query weight matrix $\mathbf{W_{q}}$ is initialized randomly, and the key weight matrix is set equal to it, $\mathbf{W_{k}}=\mathbf{W_{q}}$. This ensures that the key and query matrices are identical at initialization, introducing symmetry into the attention mechanism.